In an X – ray tube, electrons accelerated through a potential difference of 15, 000 volts strike a copper target. The speed of the emitted X – ray inside the tube is . . . . . . . . . . .m/s

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
The Bohr’s radius for nth orbit
= (n2 / z εR) r0 = 32 / 15 x 12 x 0.529 A = 3.81 A